OpenStudy (bruhitsakitkat):

how do you solve 2x+3y+6 and 2x+3y=-6?

OpenStudy (retireed):

Do you realize it is the same equation?

OpenStudy (retireed):

2x+3y+6 and 2x+3y=-6 2x +3y=-6 and 2x+3y = -6

OpenStudy (vuriffy):

When you add 6 to the other side, it is the same equation. 2x + 3y + 6 = 0.

OpenStudy (bruhitsakitkat):

no they create parallel lines one + and the other -

OpenStudy (retireed):

Nope.

OpenStudy (retireed):

Did you forget an equal sign in the first equation?

OpenStudy (falconmaster):

it is comuitive propriety

OpenStudy (vuriffy):

Graphed both, they are the same line.

OpenStudy (retireed):

I bet you had typo with the = and + key.

OpenStudy (retireed):

You might had hit shift = and got a +.

OpenStudy (bruhitsakitkat):

yes it is =6

OpenStudy (retireed):

They are parallel lines then. They newer touch. So there is no answer to the system.

OpenStudy (bruhitsakitkat):

plots (?,?)

OpenStudy (retireed):

Lines have to intersection to have a solution.

OpenStudy (vuriffy):

Your answer may be: no solution.

OpenStudy (retireed):

Sorry I don't understand plot (?, ?). There is no point to put in there.